# MediRound: Multi-Round Entity-Level Reasoning Segmentation in Medical Images

### v2

@@ Abstract
Despite the progress in medical image segmentation, most existing methods remain task-specific and lack interactivity. Although recent text-prompt-based segmentation approaches enhance user-driven and reasoning-based segmentation, they remain confined to single-round dialogues and fail to perform multi-round reasoning. In this work, we introduce Multi-Round Entity-Level Medical Reasoning Segmentation (MEMR-Seg), a new task that requires generating segmentation masks through multi-round queries with entity-level reasoning. To support this task, we construct MR-MedSeg, a large-scale dataset of 177K multi-round medical segmentation dialogues, featuring entity-based reasoning across rounds. Furthermore, we propose MediRound, an effective baseline model designed for multi-round medical reasoning segmentation. To mitigate the inherent error propagation in the chain-like pipeline of multi-round segmentation, we introduce a lightweight yet effective Judgment & Correction Mechanism during model inference. Experimental results demonstrate that our method effectively tackles the MEMR-Seg task, surpassing conventional medical referring segmentation approaches. The project is available at https://github.com/Edisonhimself/MediRound.
Figure 1: A demo dialogue of our proposed MediRound. Our model can comprehend user queries that refer to the mask results from previous rounds (e.g., the Round 2 query refers to the Round 1 mask result), enabling cross-round entity-level reasoning in multi-round medical conversations. In contrast, conventional text-prompt-based medical segmentation methods struggle in this complex task.
@@ 1 Introduction
Medical image segmentation seeks to precisely delineate regions of interest, such as organs, lesions, and tissues, across diverse medical imaging modalities [41, 31, 28]. This task supports a wide range of clinical applications [43, 15, 45, 4] and represents a fundamental enabler for continuous advancements in medical research [27]. Most existing studies focus predominantly on task-specific segmentation [26, 9, 34], commonly referred to as “specialist models”. Despite their remarkable performance, these models generally suffer from limited adaptability and lack the interactive capabilities necessary for practical, real-world deployment.
Recent advancements in text-prompt-based medical image segmentation [50, 6, 11, 20] overcome this limitation by allowing users to actively guide the model via simple textual queries, thereby enabling more diverse and user-driven segmentation outcomes. Building upon these developments, several MLLM-based approaches [39, 46, 13] further enhance the interactivity and practicality by enabling reasoning-driven medical segmentation, allowing users to obtain the desired masks through implicit queries that do not necessarily contain specialized medical terminology. However, these models primarily focus on single-round medical segmentation and are incapable of supporting multi-round and continuous text-based interactions. Moreover, although they can effectively infer implicit medical queries within a single dialogue round, they fail to perform cross-round, entity-level reasoning in multi-round conversations. As illustrated in Figure 1, in practical medical segmentation scenarios, users tend to formulate new segmentation queries based on the mask results obtained from previous rounds. For example, in Conversation 2, the user’s query is based on the mask generated in Conversation 1, and the results of Conversation 2 in turn serve as a reference for Conversation 4. Such iterative questioning not only ensures the continuity and integrity of the entire conversation but also enables users to express their desired segmentation targets more efficiently, accurately, and conveniently. Unfortunately, due to the lack of multi-round entity-level reasoning capabilities, existing methods often fail to produce the expected responses when dealing with inputs that involve cross-round logical dependencies.
@@
Given that data scarcity represents a major bottleneck in accomplishing MEMR-Seg, we first introduce a Multi-Round Entity-Level Medical Reasoning Segmentation dataset, termed MR-MedSeg, which contains a large number of medical conversations centered on entity-level reasoning segmentation with multi-round queries and answers. We collect all metadata from the publicly available SA-Med2D-20M [48]. Subsequently, inspired by SegLLM [42], we construct multi-round question–answer pairs through a combination of manual annotation and GPT-5–based generation. As a result, the MR-MedSeg dataset comprises a large and diverse collection of 177K conversations. Notably, a key distinction between the proposed MR-MedSeg and traditional medical segmentation datasets [2, 37, 10, 35, 36, 39] lies in the fact that our dataset not only features multi-round dialogues but also incorporates entity-based reasoning logic across different dialogue rounds. This design enhances the practical utility of MR-MedSeg and facilitates more interactive medical image segmentation.
Furthermore, building upon existing reasoning-based segmentation works [40, 18, 47, 42], we introduce MediRound, an effective baseline model designed for multi-round medical reasoning segmentation. MediRound integrates the segmentation feature representations from the reference round and the textual information from previous dialogues with the current query. These are jointly embedded into the input sequence of the inner-MLLM. This design enables the model not only to comprehend the current query but also to access the core information from the reference round and maintain awareness of the entire conversational history. Moreover, to address the inevitable issue of error accumulation inherent in multi-round segmentation (as illustrated in Figure 1, if the result of round 1 contains errors, these errors may propagate to round 2, and the resulting inaccuracies in round 2 can further affect round 4), we introduce a lightweight yet effective Judgment and Correction Mechanism during the model inference stage. Applied after the end-to-end model training phase, this mechanism helps MediRound mitigate error propagation in practical multi-round inference and effectively enhances the overall accuracy of multi-turn entity-level reasoning. Extensive experimental results demonstrate the superior performance of MediRound in both multi-round and single-round medical segmentation, while confirming the effectiveness of the Judgment & Correction Mechanism. Overall, we summarize our main contributions as follows:
@@
  • We propose an effective baseline model, MediRound, and introduce a Judgment & Correction Mechanism to address error accumulation during multi-round reasoning segmentation. Extensive experiments demonstrate the superior performance of our model and the utility of the proposed mechanism.
@@ 3 MR-MedSeg
Dataset Overview.   The overview of our proposed MR-MedSeg dataset is shown in Figure 2. As illustrated in the figure, we design five representative types of interaction scenarios tailored for multi-round entity-level medical reasoning segmentation, each reflecting realistic use cases in clinical practice. These scenarios each involve distinct forms of inter-instance relationships, including organ–lesion dependency, anatomical hierarchy, organ/tissue attribute association, spatial relationship, and strong inferential reasoning relationship. Figure 2 further presents the overall composition and proportions of these scenarios in MR-MedSeg. Additional detailed descriptions of each scenario and their corresponding inter-instance characteristics are provided in the Appendix.
@@ A.1 Training Pipeline
As shown in Figure A, LLaVA-Med takes $x_{all}$ (denoting the full set of inputs) and produces the [SEG] token as output. The hidden feature of [SEG], $h_{c}$, is first fed into the Correction Module $MLP_{C}$, producing $h^{\prime}_{c}$. The mask decoder $\mathcal{G}_{i}^{dec}$ then takes $h^{\prime}_{c}$ along with the visual features $f$ (extracted by the vision backbone $\mathcal{G}_{i}^{enc}$) and outputs the predicted mask $\hat{\mathbf{M}}$. The process can be formulated as:
$\begin{aligned} h^{\prime}_{c}=&\;{MLP}_{C}(h_{c}),\quad\hat{\mathbf{M}}=\mathcal{G}_{i}^{dec}(f,h^{\prime}_{c}).\end{aligned}$ (1)
@@
We leverage the end-to-end trained MediRound weights to optimize the Correction Module $MLP_{C}$ and the Quality Judgment Module $MLP_{J}$ within the JCM. During the training of JCM, all parameters of the MediRound modules (e.g., LLaVA-Med, MedSAM, etc.) are frozen, and only the Correction Module $MLP_{C}$ and the Quality Judgment Module $MLP_{J}$ are trainable.
@@ A.2 Optimization Strategy
As illustrated in Figure A, the Correction Module $MLP_{C}$ and the Quality Judgment Module $MLP_{J}$ are jointly optimized using the mask loss $\mathcal{L}_{mask}$ and the quality score loss $\mathcal{L}_{score}$.
For the quality score loss $\mathcal{L}_{score}$, we compute the Binary Cross-Entropy loss $\mathcal{L}_{bce}^{QJ}$ between the predicted quality score $q\in[0,1]$ and the ground-truth quality score $\mathbf{Q}\in[0,1]$. The ground-truth quality score $\mathbf{Q}$ is obtained by computing the Intersection-over-Union (IoU) between the predicted mask $\hat{\mathbf{M}}$ and the ground-truth mask $\mathbf{M}$, serving as a soft target in the Binary Cross-Entropy loss.
The mask loss $\mathcal{L}_{mask}$ is calculated between the predicted mask $\hat{\mathbf{M}}$ and the ground-truth mask $\mathbf{M}$. And the mask loss $\mathcal{L}_{mask}$ is a weighted sum of the per-pixel Binary Cross-Entropy $\mathcal{L}_{bce}$ and the DICE loss $\mathcal{L}_{dice}$, with the weights determined by $\lambda_{bce}$ and $\lambda_{dice}$.
The overall loss $\mathcal{L}_{JCM}$ is formulated as:
@@ E.1 Additional Interactive Cases
E.2 Data Samples of MR-MedSeg
Data samples from the MR-MedSeg dataset are presented in Figure D to illustrate the metadata structure and content. Our proposed dataset (MR-MedSeg) and model (MediRound) will all be released in the final submission version.
Figure D: Sample data in MR-MedSeg.

### v3

@@ Abstract
Despite recent progress in text-prompt-based medical image segmentation, these methods are limited to single-round dialogues and fail to support multi-round reasoning, which is important for medical education scenarios. In this work, we introduce Multi-Round Entity-Level Medical Reasoning Segmentation (MEMR-Seg), a new task that requires generating segmentation masks through multi-round queries with entity-level reasoning, helping learners progressively develop their understanding of medical knowledge. To support this task, we construct MR-MedSeg, a large-scale dataset of 177K multi-round medical segmentation dialogues, featuring entity-based reasoning across rounds. Furthermore, we propose MediRound, an effective baseline model designed for multi-round medical reasoning segmentation. To mitigate the inherent error propagation within the chain-like pipeline of multi-round segmentation, we introduce a lightweight yet effective Judgment & Correction Mechanism during model inference. Experimental results demonstrate that our method effectively addresses the MEMR-Seg task and outperforms conventional medical referring segmentation methods. The project is available at https://github.com/Edisonhimself/MediRound.
Figure 1: A medical education dialogue illustrating the proposed MediRound. Our model can comprehend user queries that refer to the mask results from previous rounds (e.g., the Round 2 query refers to the Round 1 mask result), enabling cross-round entity-level reasoning in multi-round medical conversations. This iterative querying paradigm allows learners to progressively develop their understanding of medical entities. In contrast, conventional text-prompt-based medical segmentation methods struggle with such a complex and context-dependent task.
@@ 1 Introduction
Medical image segmentation seeks to precisely delineate regions of interest (e.g., organs and lesions) across diverse medical imaging modalities [41, 31, 28]. This capability underpins a broad range of downstream medical applications, including clinical analysis, medical education, and biomedical research  [43, 15, 45, 4, 27]. Most existing studies focus predominantly on task-specific segmentation [26, 9, 34], commonly referred to as “specialist models”. Despite their remarkable performance, these models generally suffer from limited adaptability and lack the interactive capabilities necessary for practical, real-world usage.
Recent advancements in text-prompt-based medical image segmentation [50, 6, 11, 20] overcome this limitation by allowing users to actively guide the model via simple textual queries, thereby enabling more diverse and user-driven segmentation outcomes. Building upon these developments, several MLLM-based approaches [39, 46, 13] further enhance the interactivity and practicality by enabling reasoning-driven medical segmentation, allowing users to obtain the desired masks through implicit queries.
However, these models primarily focus on single-round medical segmentation and are incapable of supporting multi-round and continuous text-based interactions. Although the one-shot querying paradigm can be efficient in clinical use, it presumes high-quality prompts and substantial user expertise. As a result, it is well suited to medical experts but less conducive to the learning process of non-expert users. In medical education settings, visual understanding is typically developed progressively. As illustrated in Figure 1, learners tend to formulate new segmentation queries based on the mask results obtained from previous rounds. For example, in Round 2 Conversation, the user’s query is based on the mask generated in Round 1 Conversation, and the results of Round 2 Conversation in turn serve as a reference for Round 4 Conversation. Such iterative questioning helps learners progressively develop a better visual understanding of different medical entities and grasp the relationships among them more effectively. Moreover, these follow-up queries are often relation-driven and logically constrained, reflecting how trainees learn anatomy and pathology through relational concepts rather than precise standalone descriptions, which aligns with how learning occurs in real-world medical education. Unfortunately, due to the lack of multi-round entity-level reasoning capabilities, existing methods often fail to produce the expected responses when dealing with inputs that involve cross-round logical dependencies.
@@
Given that data scarcity represents a major bottleneck in accomplishing MEMR-Seg, we first introduce a Multi-Round Entity-Level Medical Reasoning Segmentation dataset, termed MR-MedSeg, which contains a large number of medical conversations centered on entity-level reasoning segmentation with multi-round queries and answers. We collect all metadata from the publicly available SA-Med2D-20M [48]. Subsequently, inspired by SegLLM [42], we construct multi-round question–answer pairs through a combination of manual annotation and GPT-5-based generation. As a result, the MR-MedSeg dataset comprises a large and diverse collection of 177K conversations. Notably, a key distinction between the proposed MR-MedSeg and traditional medical segmentation datasets [2, 37, 10, 35, 36, 39] lies in the fact that our dataset not only features multi-round dialogues but also incorporates entity-based reasoning logic across different dialogue rounds. This design enhances the educational utility of MR-MedSeg and facilitates more interactive medical image segmentation.
Furthermore, building upon existing reasoning-based segmentation works [40, 18, 47, 42], we introduce MediRound, an effective baseline model designed for multi-round medical reasoning segmentation. MediRound integrates the segmentation feature representations from the reference round and the textual information from previous dialogues with the current query. These are jointly embedded into the input sequence of the inner-MLLM. This design enables the model not only to comprehend the current query but also to access the core information from the reference round while maintaining awareness of the entire conversational history. Moreover, to address the inevitable issue of error accumulation inherent in multi-round segmentation (as illustrated in Figure 1, if the result of Round 1 contains errors, these errors may propagate to Round 2, and the resulting inaccuracies in Round 2 can further affect Round 4), we introduce a lightweight yet effective Judgment & Correction Mechanism during the model inference stage. Applied after the end-to-end model training phase, this mechanism helps MediRound mitigate error propagation in practical multi-round inference and effectively enhances the overall accuracy of multi-turn entity-level reasoning. Extensive experimental results demonstrate the superior performance of MediRound in both multi-round and single-round medical segmentation, while confirming the effectiveness of the Judgment & Correction Mechanism. Overall, we summarize our main contributions as follows:
@@
  • We propose an effective baseline model, MediRound, and introduce a Judgment & Correction Mechanism to address the inherent error accumulation during multi-round reasoning segmentation. Extensive experiments demonstrate the superior performance of our model and the utility of the proposed mechanism.
@@ 3 MR-MedSeg
Dataset Overview.   The overview of our proposed MR-MedSeg dataset is shown in Figure 2. As illustrated in the figure, we design five representative types of interaction scenarios tailored for multi-round entity-level medical reasoning segmentation, each reflecting realistic use cases in medical education practice. These scenarios each involve distinct forms of inter-instance relationships, including organ–lesion dependency, anatomical hierarchy, organ/tissue attribute association, spatial relationship, and strong inferential relationship. Figure 2 further presents the overall composition and proportions of these scenarios in MR-MedSeg. Additional detailed descriptions of each scenario and their corresponding inter-instance characteristics are provided in the Appendix.
@@ A.1 Training Pipeline
The overall training pipeline of JCM is illustrated in Figure A. As shown in Figure A, LLaVA-Med takes $x_{all}$ (denoting the full set of inputs) and produces the [SEG] token as output. The hidden feature of [SEG], $h_{c}$, is first fed into the Correction Module $MLP_{C}$, producing $h^{\prime}_{c}$. The mask decoder $\mathcal{G}_{i}^{dec}$ then takes $h^{\prime}_{c}$ along with the visual features $f$ (extracted by the vision backbone $\mathcal{G}_{i}^{enc}$) and outputs the predicted mask $\hat{\mathbf{M}}$. In parallel, $\mathcal{G}_{i}^{dec}$ takes $h_{c}$, $f$ and outputs the uncorrected mask $\tilde{\mathbf{M}}$. The process can be formulated as:
$h^{\prime}_{c}={MLP}_{C}(h_{c}),\quad\hat{\mathbf{M}}=\mathcal{G}_{i}^{dec}(f,h^{\prime}_{c}),\quad\tilde{\mathbf{M}}=\mathcal{G}_{i}^{dec}(f,h_{c}).$ (1)
@@
We utilize the end-to-end trained MediRound weights to optimize the Correction Module $MLP_{C}$ and the Quality Judgment Module $MLP_{J}$ within the JCM. During the training of JCM, all parameters of the MediRound modules (e.g., LLaVA-Med, MedSAM, etc.) are frozen, and only the Correction Module $MLP_{C}$ and the Quality Judgment Module $MLP_{J}$ are trainable.
@@ A.2 Optimization Strategy
For optimization, as illustrated in Figure A, the Correction Module $MLP_{C}$ and the Quality Judgment Module $MLP_{J}$ are jointly optimized using the mask loss $\mathcal{L}_{mask}$ and the quality score loss $\mathcal{L}_{score}$. For the quality score loss $\mathcal{L}_{score}$, we compute the Binary Cross-Entropy loss $\mathcal{L}_{bce}^{QJ}$ between the predicted quality score $q\in[0,1]$ and the ground-truth quality score $\mathbf{Q}\in[0,1]$. The ground-truth quality score $\mathbf{Q}$ is obtained by computing the Intersection-over-Union (IoU) between the uncorrected mask $\tilde{\mathbf{M}}$ and the ground-truth mask $\mathbf{M}$, serving as a soft target in the Binary Cross-Entropy loss $\mathcal{L}_{bce}^{QJ}$. The mask loss $\mathcal{L}_{mask}$ is also calculated between the predicted mask $\hat{\mathbf{M}}$ and the ground-truth mask $\mathbf{M}$, and it follows the same formulation as that used during MediRound training. The overall loss $\mathcal{L}_{JCM}$ can be formulated as: